# From negative to positive cosmological constant through decreasing temperature of the Universe: connection with string theory and spacetime foliation results

### v2

@@ 1 Introduction
Additional attempts achieving the anti-de Sitter (AdS) – de Sitter (dS) transition include considering two interacting dark energy fluids [5], taking running Barrow entropy into account [6], or examining quintessence fields with a negative cosmological constant [7].
@@
Solution for the sign-problem of the cosmological constant has also been suggested in the framework of quantum field theory (QFT) using a functional Renormalization Group (RG) method [8]. The concept was to assume string theory at the ultraviolet (UV) and Asymptotically Safe (AS) quantum gravity & matter at the infrared (IR) scaling regimes connected by the RG scale $k$ which serves as a bridge between the negative UV and the positive IR values of the cosmological constant. However, the RG scale $k$ is introduced artificially in the Wilsonian approach [9] and the quantized theory must be obtained in the physical limit $k\to 0$, so its use is not fully justified to connect early and late time cosmologies. In addition, temperature, which is a relevant parameter in cosmology, is missing in this approach.
@@
Thus, here we suggest to use the temperature instead of the RG scale $k$ to connect string theory at very high temperature and AS quantum gravity at intermediate and low temperatures of the expanding Universe. To achieve this one has to extend the zero-temperature functional RG method [10] to finite temperature. This can be done by using the standard finite temperature extension of QFT where the time-like dimension is compactified [11, 12, 13] and its inverse radius plays the role of the dimensionful temperature parameter $T$ (in natural units). However, the usual choice of the perturbative RG approach, when the temperature is linked to the running (perturbative) RG scale $\mu$, i.e., $T=\mu/(2\pi)$, cannot be used in the functional RG method because one has to take the limit $k\to 0$ to obtain the quantized theory. Therefore, in [14] the temperature $T$ is kept constant over the RG flow, i.e., it is linked to a fixed (not running) momentum scale. Most of the further literature on thermal functional RG is based on this assumption. In this fixed $T$ approach, the dimensionful temperature is well defined and the limit $k\to 0$ can be done safely, but there is a price to pay: the RG flow equations have no fixed point solutions. Therefore, progress in this direction has been hampered by the fact that the usual extension of thermal functional RG method at finite temperature does not reproduce non-trivial fixed points needed for the RG analysis of various questions formulated in the framework of thermal QFT.
To overcome these difficulties, in Refs. [15, 16, 17] we proposed a modification of the usual functional RG approach af finite temperature by relating the temperature parameter to the running RG scale as $T\equiv k_{T}=\tau k$ (in natural units). In our approach
$\tau=T/k$
@@
is kept constant over the RG flow while taking the simultaneous $T,k\to 0$ limit. Thus the dimensionful parameter $T$ changes by the RG scale, so it is not considered as the temperature, but rather understood as a running cutoff for thermal fluctuation. The dimensionless $\tau$ is identified as the physical temperature of the surrounding plasma. In this case the RG flow equations have real (and not pseudo) fixed points, which have fundamental importance since the determination of critical behaviour and various phases are strongly related to fixed point solutions of RG flow equations. In summary, in the fixed $T$ approach the dimensionful temperature is well defined, but RG flow equations have no fixed points. In our fixed $\tau$ approach the dimensionful temperature is not defined, but the fixed points are. This small, but crucial modification of the original thermal RG method opened the avenue to consider the interplay of classical (CPT) and quantum (QPT) phase transitions, as discussed in [17] which gave us the possibility to compare it to simulation results and to confirm the viability of the fixed $\tau$ scheme. Moreover, in the recent work [16] it was studied the thermal RG study of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ approach. The quantum effective action at a given dimensionless temperature $\tau$ was given by moving along the thermal RG trajectory and this procedure was repeated for various values of $\tau$, which resulted in the $\tau$-evolution of the Reuter [18] (i.e., non-Gaussian UV) fixed point. We showed that in the high temperature limit ($\tau\to\infty$) the $g$-coordinate of the Reuter fixed point vanishes and the cosmological constant takes on a negative value in the limit $k\to 0$. It is not in disagreement with observations, since during the thermal evolution of the Universe a thermal phase transition occurs and the cosmological constant runs to the expected positive value at low temperatures. This mechanism to solve the sign-problem of the cosmological constant is represented on Fig. 1.
@@
Motivated by the results in [16], here we study the generality of these findings. We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [19, 20]) has the cosmological constant with a negative (positive, respectively) sign for large (small) temperature. Therefore, the scheme represented in Fig. 1 can be used in every case which can be seen as a natural solution for the sign problem of the cosmological constant. In particular, we study three extensions/variants of the simplest AS quantum gravity: the conformally reduced and the ghost-improved versions and its extension by scalar matter fields. In addition to that we discuss the connection to spacetime foliation.
@@ 4 Ghost-improved EH gravity at $T=0$
where $C,\bar{C}$ are the classical ghost fields, and $c,\bar{c}$ are their associated background fields: one has $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$, $\bar{C}_{\mu}=\bar{c}_{\mu}+\bar{f}_{\mu}$ and $C_{\mu}=c_{\mu}+f_{\mu}$, where $h_{\mu\nu}$, $f_{\mu}$, $\bar{f}_{\mu}$ mark the expectation value of the quantum fluctuations around the background. The ansatz is constructed from the $\Gamma_{k}^{\text{grav}}$ gravitational term, the $\Gamma_{k}^{\text{gf}}$ gauge-fixing term and the $\Gamma_{k}^{\text{gh}}$ ghost term, with their explicit forms also given in [24]. The harmonic gauge choice is also used in order to compare the result to the EH truncated gravity without ghost-improvement. Beta-functions in $d$ dimensions are found to be
@@ 5 AS gravity at finite temperature
Figure 2: Thermal RG flow diagrams of CREH, QEG with $N$-component scalar field (with chosen parameters $\xi=10$ and $N=2$) and Ghost-improved EH gravity for various $\tau$ dimensionless temperature values. With $\tau\to\infty$ the $g^{*}\to 0$ limit is reached in each case.
@@ Acknowledgements
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” are gratefully acknowledged. The creation of this scientific communication was supported by the University of Miskolc with funding granted to the author I. Nándori within the framework of the institution’s Scientific Excellence Support Program. (Project identifier: ME-TKTP-2025-3). This work was supported by the University Research Scholarship Programme (EKÖP-25-2-DE-376) of the Ministry of Culture and Innovation and of the National Research, Development and Innovation Fund.
@@ Appendix A Functional RG study of AS gravity at $T=0$
One can recover the Einstein-Hilbert truncated anomalous dimension without ghost-improvements by setting $C_{i}(\lambda)=0$.
@@ Appendix B Arnowitt-Deser-Misner formalism
We summarise here the main ideas behind the Arnowitt-Deser-Misner (ADM) formalism used in the discussion on the connection between spacetime foliation and thermal RG approach. In general relativity the metric is equipped with Lorentzian signature, however $\Gamma_{k}$ is given by an Euclidean path integral therefore the Lorentzian spacetime has to be recovered by Wick-rotation. QFT calculations are usually done on a fixed Minkowski background metric, which provides a clear notion of time. However, when discussing dynamical spacetimes, the role of time can be questioned. To address this the ADM-formalism is used, in which the spacetime metric $g_{\mu\nu}$ is decomposed into a lapse function $N_{l}$, a shift-vector $N_{i}$ and a metric on spatial slices $\sigma_{ij}$. These are needed, because the $4D$ spacetime is sliced into $3D$ spatial hypersurfaces with metric $\sigma_{ij}$, each hypersurface labeled with time parameter $t$. $N_{l}$ – being related to the separation between hypersurfaces – and $N_{i}$ – which is a displacement related to a point passing to the next surface – describe how to weld the hypersurfaces together to form the foliation, therefore imprinting the Euclidean spacetime with a distinguished direction. The resulting preferred time direction enables the computation of transition amplitudes from an initial to a final slice.

### v3

@@ 1 Introduction
Additional attempts achieving the anti-de Sitter (AdS) – de Sitter (dS) transition include considering two interacting dark energy fluids [6], taking running Barrow entropy into account [7], or examining quintessence fields with a negative cosmological constant [8]. Another example, for the AdS–dS transition is given in [9] where the generalization of the Hawking-Page phase transition [10] is used.
The hypothetical AdS–dS transition – which is said to alleviate the $H_{0}$ tension [11, 12, 13] and the $S_{8}$ discrepancy [14, 15, 16] – is strongly influenced by the AdS Swampland conjecture [17, 18, 19] because the logarithmic dependence ($\log|\Lambda|$) makes it impossible to cross the $\Lambda=0$ barrier at zero temperature [20, 21]. However, it was argued and shown in [20] that at finite temperature the number of light particles can change making the sign switch possible.
[…]
@@
Solution for the sign-problem of the cosmological constant has also been suggested in the framework of quantum field theory (QFT) using a Renormalization Group (RG) in particular, the Functional Renormalization Group (FRG) method [28]. The concept was to assume string theory at the ultraviolet (UV) and Asymptotically Safe (AS) quantum gravity & matter at the infrared (IR) scaling regimes connected by the RG scale $k$ which serves as a bridge between the negative UV and the positive IR values of the cosmological constant. However, the RG scale $k$ is introduced artificially in the Wilsonian approach [29] and the quantized theory must be obtained in the physical limit $k\to 0$, so its use is not fully justified to connect early and late time cosmologies. In addition, temperature, which is a relevant parameter in cosmology, is missing in this approach.
@@
Thus, here we suggest to use the temperature instead of the RG scale $k$ to connect string theory at very high temperature and AS quantum gravity at intermediate and low temperatures of the expanding Universe. To achieve this one has to extend the zero-temperature FRG method [30] to finite temperature. This can be done by using the standard finite temperature extension of QFT where the time-like dimension is compactified [31, 32, 33] and its inverse radius plays the role of the dimensionful temperature parameter $T$ (in natural units).
In the usual finite-temperature perturbative RG approach the temperature parameter $T$ and the running (perturbative) RG scale $\mu$ are linked to each other $\mu=2\pi T$. The RG scales of the perturbative and the non-perturbative approaches are connected $\mu\sim k$. However, the $k$-dependence of the effective action of the non-perturbative FRG approach is introduced artificially to implement the Wilsonian integration of fluctuating modes and the quantised theory is obtained in the physical limit $k\to 0$ which has to be taken unavoidably. Thus, in the usual finite temperature FRG approach, the dimensionful temperature parameter $T$ is linked to a fixed (not running) momentum scale $\Lambda$ which results in $T=\tau\Lambda$, see [34]. Most of the further literature on thermal FRG is based on this assumption. In this fixed $T$ approach, the dimensionful temperature is well defined and the limit $k\to 0$ can be done safely, but there is a price to pay: the RG flow equations have no non-trivial fixed point solutions. To overcome these difficulties, in Refs. [35, 36, 37] a modification of the usual FRG approach at finite temperature was proposed by relating the temperature parameter to the running RG scale as $T=\tau k$ (in natural units). In our approach
$\tau=T/k$ (1)
is kept constant over the RG flow while taking the simultaneous $T,k\to 0$ limit. Thus the dimensionful parameter $T$ changes by the RG scale.
Our working hypothesis is then to consider the dimensionful temperature parameter $T$ as a running cutoff for thermal fluctuation and to fix the parameter $\tau$ over the RG flow according (1): $T=\tau k$, where $T$ no longer plays the role of the temperature. We will use the notation $T\equiv k_{T}$ where $k_{T}$ is understood as the running cutoff for thermal and $k$ is for quantum fluctuations. In our approach the “true” temperature (the one entering the thermodynamic quantities such as the free energy that one aims at calculating) is proportional to the dimensionless quantity $\tau$ which is kept constant over the RG flow; it can be chosen arbitrarily, it is not restricted to a specific value and therefore can have any finite value. The relation between $\tau$ and the “true” temperature $T_{true}$ is $\tau=k_{B}T_{true}/J$, where $J$ is an energy scale of the system depending on the microscopic details of the system. Since our results depend only on $\tau$ and not separately on $T_{true}$ and $J$, we keep the discussion in terms of $\tau$. In other words, one can say $\tau$ measures how large the thermal fluctuations are compared to the quantum ones. Thus, our thermal RG scheme differs from the usual non-perturbative and also from the perturbative ones. The rest of the paper is devoted to understand the consequences of our working hypothesis on the thermal evolution of the cosmological constant.
@@
In summary, in the fixed $T$ approach the dimensionful temperature is well defined, but RG flow equations have no fixed points. In our fixed $\tau$ approach the dimensionful temperature is not defined, but the fixed points are. This small, but crucial modification of the original thermal RG method opened the avenue to consider the interplay of classical (CPT) and quantum (QPT) phase transitions, as discussed in [37] which gave us the possibility to compare it to simulation results and to confirm the viability of the fixed $\tau$ scheme. Moreover, in the recent work [36] we studied the thermal RG of the simplest AS quantum gravity model, the Einstein-Hilbert (EH) truncation, using the fixed $\tau$ approach. The quantum effective action at a given dimensionless temperature $\tau$ was given by moving along the thermal RG trajectory and this procedure was repeated for various values of $\tau$, which resulted in the $\tau$-evolution of the Reuter [38] (i.e., non-Gaussian UV) fixed point. We showed that in the high temperature limit ($\tau\to\infty$) the $g$-coordinate of the Reuter fixed point vanishes and the cosmological constant takes on a negative value in the limit $k\to 0$. It is not in disagreement with observations, since during the thermal evolution of the Universe a thermal phase transition occurs and the cosmological constant runs to the expected positive value at low temperatures. This mechanism to solve the sign-problem of the cosmological constant is represented on Fig. 1.
@@
Motivated by the results of [36], here we study the generality of these findings. We find that the thermal RG study of basically any AS quantum gravity model (for recent reviews see [39, 40]) has the cosmological constant with a negative (positive, respectively) sign for large (small) temperature. Therefore, the scheme represented in Fig. 1 can be used in every case which can be seen as a natural solution for the sign problem of the cosmological constant. In particular, we study three extensions/variants of the simplest AS quantum gravity: the conformally reduced and the ghost-improved versions and its extension by scalar matter fields. In addition to that we discuss the connection to spacetime foliation.
@@ 4 Ghost-improved EH gravity at $T=0$
where $C,\bar{C}$ are the classical ghost fields, and $c,\bar{c}$ are their associated background fields: one has $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$, $\bar{C}_{\mu}=\bar{c}_{\mu}+\bar{f}_{\mu}$ and $C_{\mu}=c_{\mu}+f_{\mu}$, where $h_{\mu\nu}$, $f_{\mu}$, $\bar{f}_{\mu}$ mark the expectation value of the quantum fluctuations around the background. The ansatz is constructed from the $\Gamma_{k}^{\text{grav}}$ gravitational term, the $\Gamma_{k}^{\text{gf}}$ gauge-fixing term and the $\Gamma_{k}^{\text{gh}}$ ghost term, with their explicit forms also given in [44]. The harmonic gauge choice is also used in order to compare the result to the EH truncated gravity without ghost-improvement. Beta-functions in $d$ dimensions are
@@ 5 AS gravity at finite temperature
Figure 2: Thermal RG flow diagrams of CREH, QEG with $N$-component scalar field (with $\xi=10$ and $N=2$) and Ghost-improved EH gravity for various $\tau$ dimensionless temperature values. With $\tau\to\infty$ the $g^{*}\to 0$ limit is reached in each case.
@@ Acknowledgements
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” are gratefully acknowledged. This work was supported by the University Research Scholarship Programme (EKÖP-25-2-DE-376) of the Ministry of Culture and Innovation and of the National Research, Development and Innovation Fund.
Note added
After the submission several interesting papers appeared [26, 27, 57, 58, 59, 60, 61], where the authors discuss observational results in terms of a phenomenological model supporting a scenario compatible with our results. Further investigation on this point is currently on-going.
@@ Appendix A Functional RG study of AS gravity at $T=0$
One can recover the Einstein-Hilbert truncated anomalous dimension without ghost-improvements by setting $C_{i}=0$.
@@ Appendix B Arnowitt-Deser-Misner formalism
The Arnowitt-Deser-Misner (ADM) formalism is used to connect spacetime foliation and thermal RG. In general relativity the metric is equipped with Lorentzian signature, however $\Gamma_{k}$ is given by an Euclidean path integral thus the Lorentzian spacetime has to be recovered by Wick-rotation. QFT calculations are usually done on a fixed Minkowski background which provides a clear notion of time. However, when discussing dynamical spacetimes, the role of time can be questioned. To address this the ADM-formalism is used, in which the spacetime metric $g_{\mu\nu}$ is decomposed into a lapse function $N_{l}$, a shift-vector $N_{i}$ and a metric on spatial slices $\sigma_{ij}$. These are needed, because the $4D$ spacetime is sliced into $3D$ spatial hypersurfaces with metric $\sigma_{ij}$, each hypersurface labeled with time parameter $t$. $N_{l}$ – being related to the separation between hypersurfaces – and $N_{i}$ – which is a displacement related to a point passing to the next surface – describe how to weld the hypersurfaces together to form the foliation, thus imprinting the Euclidean spacetime with a distinguished direction. The resulting preferred time direction enables the computation of transition amplitudes from an initial to a final slice.